# Copyright Infringement Detection in Text-to-Image Diffusion Models via Differential Privacy

### v2

@@ 6 Hierarchical Categories of Copyright Infringement and CIDD Dataset
While existing datasets provide preliminary resources for copyright infringement detection, they have several limitations due to strong and unrealistic assumptions. CopyScope (Zhou et al., 2023) and CPDM (Ma et al., 2024) primarily focus on infringed examples and lack non-infringed counterparts, which makes it difficult to construct reliable image-label pairs or evaluate detection thresholds. Other benchmarks (Wang et al., 2024) only cover a few iconic character concepts, and fail to generalize to broader cases.
To overcome these limitations, we construct the Copyright Infringement Detection Dataset (CIDD). It contains several classes of orthogonal prompts and three image classes that are most likely to be infringed: human face, architecture, and arts painting, with a total of 429 concepts and 2,397 images. Each class is mapped to one hierarchical category in table 1. Crucially, CIDD includes both infringed and non-infringed concepts, each of which is annotated with a binary infringement label based on its source and content provenance, and is paired with 3 to 6 neighbourhood images, enabling robust learning and evaluation under weak and probabilistic assumptions. Data examples, summaries, and collection methods are detailed in Appendix B.
7 Experiment
  Class SD1.4 SDXL-1.0 SANA-0.6B FLUX.1
Human Face 0.9011 0.8058 0.7011 0.6289 0.8062 0.7285 0.7531 0.6419
Merged Total 0.8071 0.6726 0.7800 0.7234 0.7914 0.6855 0.8257 0.7039
As mentioned in Section 2, Copyscope cannot distinguish image-level infringement; DIAGNOSIS is an infringement detection method based on a priori dataset coating; CPDM can only evaluate the extent of infringement after data unlearning. These methods target unrealistic settings and different goals, making direct comparison with DPM inappropriate. As for detection in LVLMs, DPM can also be extended to this setting, which we leave as a promising work in the future (see Appendix F).
In this section, we evaluate the proposed DPM on the CIDD dataset across four text-to-image models: Stable Diffusion v1.4 (SD1.4) (Rombach et al., 2022), Stable Diffusion XL (SDXL) (Podell et al., 2023), SANA (Xie et al., 2024), and FLUX (Labs et al., 2025) models. We report both quantitative and qualitative results, and perform an ablation study to validate all modules in DPM. For the quantitative evaluation, we employ two standard metrics: ROC-AUC, which measures the ability to distinguish infringed from non-infringed samples, and Soft Accuracy (SoftAcc), a stricter metric that evaluates the numerical alignment between confidence scores and ground truth labels, penalizing even slight deviations from the correct value. Details of the metrics are listed in the Appendix D.1, and experiment settings in Appendix D.2.
7.1 Quantitative Experiments
We report per-class and aggregated results in Table 2. DPM consistently achieves a weighted-average AUC above 80% and SoftAcc above 72% across models, demonstrating its strong generalization. Moreover, detection performance varies across classes. For instance, DPM achieves only 70% AUC on the human face class when using SDXL-1.0, mainly because ViT-L/14 produces more generalized embeddings, and the more diverse training data entangles human faces with common visual patterns, thereby weakening DPM’s sensitivity response.
Figure 3: Qualitative visualization of the Unlearning Branch and Learning Branch across different timesteps. Models tend to learn and unlearn faster with infringed samples, while slower on non-infringed ones, and cannot learn exact elements in the target images.
7.2 Qualitative Experiments
To better understand the behaviors of our two-branch framework, we visualize representative samples in Fig. 4. These results suggest that the infringed samples have triggered the model’s internal memory, causing it to reproduce memorized content with high fidelity; while the unlearning branch rapidly suppresses such resemblance, producing visually dissimilar outputs after just a few negative steps. For non-infringed samples, both branches tend to maintain stable generations with limited directional change. We also visualize the AUC-ROC curves in Appendix D.3.
7.3 Ablation Study and Robustness
We provide a detailed ablation study and robustness analysis (see Appendix D.4 & D.5) to examine the impact of different components in our detection framework. We find that: (1) the proposed conditional sensitivity metric effectively distinguishes between infringed and non-infringed samples, as partially illustrated in Fig. 4; (2) merging of two branches and multiple timesteps both improve and stabilize detection performance. We also discuss limitations and future work in Appendix E and F respectively.
8 Conclusion
Appendix A Challenges of Copyright Infringement Detection
Scalability of Detection
Modern training dataset may contain trillions of images, making exhaustive matching or retrieval computationally infeasible. Efficient detection methods must therefore avoid reliance on dataset-wide comparisons.
Inaccessibility of Training Data
Training datasets are often proprietary and not publicly disclosed due to legal, commercial, or privacy constraints. In some cases, even the model developers lack full visibility into the data sources and do not retain a complete record of all data sources, making direct access or auditing impossible.
Input Unavailability in Conditional Models
In many real-world scenarios, only the generated image is publicly observable (e.g., on social media), while the original corresponding input (e.g., text prompt $t$ or conditioning image $x$) is missing or irretrievable. Consequently, detection must be based solely on the generated content, without knowledge of its conditioning inputs.
Insufficient Theoretical Guarantees
Most existing detection techniques rely on heuristic or empirical signals, lacking formal theoretical guarantees. While such methods may indicate suspicious behavior, they do not constitute definitive evidence of training set inclusion, limiting their legal and forensic utility.
Appendix B CIDD Preview
@@
All images are validated for their existence by: (1) CLIP-based retrieval within the relevant pretraining datasets; (2) direct text-to-image model generation with the corresponding target concept prompts. In addition, non-infringed samples are validated on the website https://haveibeentrained.com/ to ensure their absence from any known training corpus.
@@ Appendix C Details of Hierarchical Categories in Copyright Infringement
As shown in Fig. 6, we elaborate on the five hierarchical categories of content resemblance that consolidate our framework for detecting copyright infringement in generative models. Each level reflects a different degree of abstraction, from superficial visual similarity to deep conceptual overlap, which reflects increasing levels of abstraction and creative expression:
@@
Figure 6: Hierarchical Categories in Copyright Infringement. It is ordered from low-level perceptual features to high-level conceptual constructs.
@@ D.3 Quantative and Qualitative Experiments
The relatively higher variance in the non-infringed group may be attributed to two factors: (1) some non-infringed concepts, although not memorized, may closely resemble commonly seen data in the training dataset, leading to relatively high DPM scores; (2) certain non-infringed concepts are inherently more unique or abstract, and exhibit lower sensitivity and DPM scores, which further increases score dispersion.
@@
Figure 7: ROC curves in four representative models. It shows DPM confidence score in learning branch, unlearning branch, and merged branches.
@@
Also, in Fig. 7, we observe that the SANA-0.6B model underperforms in the learning branch, showing weaker separation between infringed and non-infringed subsets compared to other models. Specifically, Fig. 9 shows that the sensitivity metric in the learning branch increases slowly, and the distributions of infringed and non-infringed subsets remain closely overlapping. We attribute this behavior to several key factors: (1) limited parameter capacity of the 0.6B-sized SANA model restricts its ability to memorize complex concepts; (2) the use of parameter-efficient tuning methods such as LoRA may further limit the expressive adaptation needed to internalize new concepts during fine-tuning. These factors jointly lead to reduced sensitivity shifts, weakening the contrastive signal upon which DPM relies.
@@ D.4 Ablation Study
Figures 9 and 11 illustrate the evolution of sensitivity scores across different model architectures and CIDD categories respectively. It shows that in most cases, the interquartile ranges of the two subsets show minimal overlap, reinforcing the discriminative power of the metric. As fine-tuning progresses, a clear and consistent gap emerges between the two subsets, with infringed samples exhibiting substantially higher sensitivity scores. This observation confirms that DPM is capable of reliably detecting memorized content and demonstrates strong generalizability across both models and semantic domains.
In Fig. 9, we observe that FLUX.1 model exhibits more outliers in the learning branch. This instability can be attributed to the relatively large learning rate applied to accelerate fine-tuning on this larger model, which is not caused by deficiencies in the proposed DPM framework itself, but stems from a tradeoff in fine-tuning strategy. However, due to the large parameter scale of FLUX.1, this aggressive setting may lead to unstable training dynamics or even divergence. As illustrated in Fig. 10, the fine-tuned model failed to generate customized image that affects the sensitivity metric.
Figure 10: Examples of failed generations in the learning branch of FLUX.1 at different training steps. vicky is a non-infringed concept, while david_beckham is an infringed one. Both collapses as the timestep increases, primarily due to the aggressive learning rate rather than any flaw in our DPM framework.
Figure 11: Conditional sensitivity metric across three CIDD categories in SD1.4: (left) Human Face, (middle) Architecture, and (right) Arts Painting.
@@
Due to the large computational cost of generating orthogonal prompts, we randomly selected about 60 samples from the human face category, and sampled from more than 200 orthogonal prompts. As shown in table 7, incorporating statistical normalization yields a slight but consistent improvement in both AUC and Soft Accuracy. This suggests that rectifying global parameter shifts enhances the robustness of the conditional sensitivity metric. While the gain in better performance is modest, the approach provides stronger statistical guarantees and greater interpretability.
@@ D.5 Robustness
To further evaluate the robustness of our method under real-world perturbations, we evaluate the DPM framework against two common types of image degradation: (1) lossy image quality compression with quality factor $IQ(\text{jpeg})=30\%$, and (2) additive Gaussian noise with standard deviation $\sigma=0.03$.
As shown in table 8, DPM maintains strong performance under both perturbation settings, with only minor fluctuations in AUC and Soft Accuracy.
@@
Appendix E Limitations
While the proposed DPM framework demonstrates strong performance across models and categories, it has several inherent limitations.
First, DPM relies on a strong conditional publicity signal, which may be weak when the learning branch fails to simulate memorization of the target concept, such as when the concept is highly entangled with common visual priors (e.g., generic architectural styles).
Furthermore, CLIP image embeddings used for quantifying conditional publicity can be refined. We observe that concepts with shared semantics or structures can produce similar representations, making it challenging to distinguish genuine memorization from general semantic understanding.
Lastly, while the CIDD dataset provides a valuable benchmark with binary ground-truth labels (infringed vs. non-infringed), future extensions could include richer annotations and more accurate measurements of infringement. For example, instead of assigning a binary label to each concept, we could quantify the relative degree of infringement by measuring the frequency of a given concept’s images within the pretraining dataset. Such statistical indicators would provide a finer-grained view of memorization risk and enable more scalable, probabilistic detection frameworks.
Appendix F Future Work

### v3

@@ Appendix B Copyright Infringement Detection Dataset (CIDD) Preview
While existing datasets provide preliminary resources for copyright infringement detection, they have several limitations due to strong and unrealistic assumptions. CopyScope (zhou2023copyscopemodellevelcopyrightinfringement) and CPDM (ma2024datasetbenchmarkcopyrightinfringement) primarily focus on infringed examples and lack non-infringed counterparts, which makes it difficult to construct reliable image-label pairs or evaluate detection thresholds. Other benchmarks (wang2024evaluatingmitigatingipinfringement) only cover a few iconic character concepts, and fail to generalize to broader cases.
Summaries
@@
Intra-Class Similarity
As shown in the table 4, Human face naturally has the highest similarity and is fine-grained, making semantic distinction subtler and narrowing the gap in conditional sensitivity.
  Class CLIP Similarity
Human Face 0.85
Architecture 0.79
Arts Painting 0.70
Table 4: CLIP similarity across three categories in CIDD dataset.
@@
All images are validated for their existence by:
(1) CLIP-based retrieval within the relevant pretraining datasets;
(2) direct text-to-image model generation with the corresponding target concept prompts.
In addition, non-infringed samples are validated on the website https://haveibeentrained.com/ to ensure their absence from any known training corpus.
@@ Appendix C Details of Hierarchical Categories in Copyright Infringement
We propose five hierarchical categories of content resemblance that consolidate our framework for detecting copyright infringement in generative models. Each level reflects a different degree of abstraction, from superficial visual similarity to deep conceptual overlap, which reflects increasing levels of abstraction and creative expression:
@@ D.3 Quantative and Qualitative Experiments
The relatively higher variance in the non-infringed group may be attributed to two factors:
(1) Some non-infringed concepts, although not memorized, may closely resemble commonly seen data in the training dataset, leading to relatively high DPM scores;
(2) Certain non-infringed concepts are inherently more unique or abstract, and exhibit lower sensitivity and DPM scores, which further increases score dispersion.
@@
Also, in Fig. 4, we observe that the SANA-0.6B model underperforms in the learning branch, showing weaker separation between infringed and non-infringed subsets compared to other models. Specifically, Fig. 8 shows that the sensitivity metric in the learning branch increases slowly, and the distributions of infringed and non-infringed subsets remain closely overlapping. We attribute this behavior to several key factors:
(1) limited parameter capacity of the 0.6B-sized SANA model restricts its ability to memorize complex concepts;
(2) the use of parameter-efficient tuning methods such as LoRA may further limit the expressive adaptation needed to internalize new concepts during fine-tuning.
These factors jointly lead to reduced sensitivity shifts, weakening the contrastive signal upon which DPM relies.
@@ D.4 Ablation Study
Figure 9: Conditional sensitivity metric across three CIDD categories in SD1.4: (left) Human Face, (middle) Architecture, and (right) Arts Painting. It shows that infringed samples are more sensitive to the model outputs behavior.
Figures 8 and 9 illustrate the evolution of sensitivity scores across different model architectures and CIDD categories respectively. It shows that in most cases, the interquartile ranges of the two subsets show minimal overlap, reinforcing the discriminative power of the metric. As fine-tuning progresses, a clear and consistent gap emerges between the two subsets, with infringed samples exhibiting substantially higher sensitivity scores. This observation confirms that DPM is capable of reliably detecting memorized content and demonstrates strong generalizability across both models and semantic domains.
@@
Due to the large computational cost of generating orthogonal prompts, we randomly selected about 60 samples from the human face category, and sampled from more than 200 orthogonal prompts.
As shown in table 8, incorporating statistical normalization yields a slight but consistent improvement in both AUC and Soft Accuracy. This suggests that rectifying global parameter shifts enhances the robustness of the conditional sensitivity metric. While the gain in better performance is modest, the approach provides stronger statistical guarantees and greater interpretability.
@@ D.5 Robustness
To further evaluate the robustness of our method under real-world perturbations, we evaluate the DPM framework against two common types of image degradation:
(1) lossy image quality compression with quality factor $IQ(\text{jpeg})=30\%$;
(2) additive Gaussian noise with standard deviation $\sigma=0.03$.
As shown in table 9, DPM maintains strong performance under both perturbation settings, with only minor fluctuations in AUC and Soft Accuracy. It is because DPM relies on the detection of deep semantic and stylistic features, rather than superficial pixel changes.
@@
Similarly, different templates affect little to model parameters and won’t affect our method:
  Prompt Templates AUC SoftAcc
a photo of [V] painting (Original) 0.86 0.76
a photo of sks painting 0.82 0.74
a painting in the style of V* art 0.87 0.79
randomized prompts 0.86 0.77
Table 10: Evaluation of DPM Confidence Scores under different prompt templates in SD1.4, arts painting.
Appendix E Future Work